23. August 2016: Dienstag

LDAP/IdM:

  • Nightly Update: Two issues
  • Accounts: No issues
  • Issues:
    • OSP name change did not work out:
      • I had updated the LDAP entry to make it work, but it collided with PeopleSoft’s saved netid list.
      • Put the NetID back to the old value and rerun did the work.
    • PS new faculty, new name:
      • Enough matches it is the same one. Set up for it to work tonight.
  • Development:
    • Updated remove_account to nuke group memberships as well.

Backups:

  • Save Groups:
    • One client was replaced yesterday and the NetWorker software not installed – fixed.
    • One client fell victim of the database problem – fixed.
  • Clones:
    • Started at 11:27
    • COMIS completed at 12:23
    • TAPE completed at 15:32
  • Tapes:
    • Reclaimed 3 tapes for the vaults.
    • Fix retention times on 13 savesets across five (5) tapes.
  • Issues:
    • Tape Drive *49c93 vs tape 110183L4:
      • Qualstar is not buying that the drive has failed this time.
      • Rebooted – back in service – watch like a hawk.
    • XLS fan0 failed:
      • Qualstar has issued RMA # 50342.
      • Fan0 is working today… but I don’t trust it.
      • New fan will arrive tomorrow.

Tags:

Leave a Reply

You must be logged in to post a comment.